OpenStudy (anonymous):

What's the solution to the system of linear equations 2x + y = 6 x - y = 6 Please show me what you did to get this answer step by step :-)

OpenStudy (bahrom7893):

From the second equation: x-y=6 add y to both sides: x=6+y. Ok?

OpenStudy (bahrom7893):

wait there's an even easier way.

OpenStudy (anonymous):

Set y= to the rest of the stuff. And then set those equations equal to eachother. If you can remember, that, then you can solve any of these problems. y=6-2x y=x-6 You try setting them equal to each other and solve for x. 6-2x=x-6

OpenStudy (bahrom7893):

2x + y = 6 x - y = 6 add the equations: 3x = 12 x = 4

OpenStudy (bahrom7893):

x-y=6 4-y=6 y = -2

OpenStudy (anonymous):

Okay I get that. Thank you :-)

OpenStudy (sasogeek):

\[\huge \text{Awesome!} \]

OpenStudy (anonymous):

So then when you find your x-value, you can plug that in to EITHER equation and you'll get the same y-value. Try it for yourself to verify :-)

OpenStudy (anonymous):

And if you get an equation that is y=2x+4 and y=2x+3, here's what happens: 2x+4=2x+3 (subtract from both sides) 0+4=3 0=-1 (zero does not equal -1, the y-values in the respective equations will never have any x-value in common). That means they're parallel since there's no intersection.

OpenStudy (anonymous):

I meant to say (subtract 2x from both sides)

OpenStudy (anonymous):

The answer bahrom gave me didn't add up. the second equation gave me a -6 instead of a 6.

OpenStudy (anonymous):

6-2x = x-6 6-3x = -6 -3x = -12 x = 4

OpenStudy (anonymous):

2x + y = 6 x - y = 6 3x=6 x=2

OpenStudy (anonymous):

ah crap, I need to do this stuff on paper instead of the computer!

OpenStudy (anonymous):

Sorry i meant 12 so x = 4

OpenStudy (anonymous):

Okay but what is y?

OpenStudy (anonymous):

brizy, look at my solution

OpenStudy (anonymous):

Just plug in x to either one of the equations

OpenStudy (anonymous):

bahrom gave me -2 as the y answer but i disagree with that.

OpenStudy (anonymous):

Now, plugging x=4 back into the first equation: 2(4)+y = 6 y = 6-8 y=-2 Plugging x=4 into the second equation: 4-y = 6 -y = 2 y = -2 Therefore, x=4 works since y is the same value.

OpenStudy (sasogeek):

i'm not sure how u got a -6 instead of a 6. but here's how it works x=4 y=-2 taking the same second equation you were having trouble to add up, here's what happens (4)-(-2)=6 2 negatives is a possitive, implying that 4+2=6 6=6

OpenStudy (anonymous):

I'm only in 8th grade math.

OpenStudy (anonymous):

Just look over what I typed in. Does that make sense when I plugged the x-value into both equations?

OpenStudy (sasogeek):

ohhh 8th grade :D awesome!, you should play with these equations :) i hope you understand now though :)

OpenStudy (anonymous):

Another thing you can play around is plot both of these by hand on paper. Look at the line of intersection. It's going to be where x=4 and y=-2. If you keep staring at it long enough, you'll have a better understanding of these lines.

OpenStudy (anonymous):

Ok I did what you said brinethery. I got it now. Thank you guys :-)

OpenStudy (anonymous):

great! That's what I wanted to hear :-)

OpenStudy (sasogeek):

anytime :)

OpenStudy (anonymous):

Brizy, now you can help me with my linear algebra ;;-)

OpenStudy (anonymous):

Lol yeah :-)